OpenStudy (anonymous):

x^4 + 3x^2 - 2 = 0 Solve for x

OpenStudy (anonymous):

Let a=x^2 then x^4 + 3x^2 - 2 = 0 (x^2)^2+3x^2-2=0 a^2+3a-2=0

Parth (parthkohli):

Let \(x^2 = y\):\[y^2 + 3y - 2 = 0 \]Solve for \(y\).

OpenStudy (anonymous):

Now, can u solve for a

OpenStudy (lgbasallote):

can i use z next?

OpenStudy (tyteen4a03):

@lgbasallote Be my guest ;)

OpenStudy (anonymous):

I know about solving this with substitution. I just got a weird number so I'm just wondering it that's right

OpenStudy (lgbasallote):

let \(z = x^2\) \[\implies z ^2 + 3z - 2 = 0\] factor out to solve for z

OpenStudy (lgbasallote):

and no..it shouldn't be a weird number

OpenStudy (lgbasallote):

on second thought..it is

OpenStudy (lgbasallote):

you have to use Q.F. and not factoring

OpenStudy (anonymous):

USe quadratic formula

Parth (parthkohli):

\[z = {-b \pm \sqrt{b^2 - 4ac} \over 2a} \]

OpenStudy (anonymous):

you need to use quadratic formula and the result is weird number....

Parth (parthkohli):

I hate weird @Numbers.

OpenStudy (lgbasallote):

her engagement got a lot down since i dethroned her....

Parth (parthkohli):

agdg's too....

OpenStudy (lgbasallote):

@rebecca1233 now has the second highest engagement

OpenStudy (lgbasallote):

wonder when she'll dethrone me >:))

OpenStudy (anonymous):

@lgbasallote funny haan!!!!

Parth (parthkohli):

Weird thing is that agd's SmartScore is still increasing.

Parth (parthkohli):

Satellite will be dethroned by @lgbasallote

OpenStudy (lgbasallote):

i have to go through the triumvirate first =_= the royal bodyguards of zeus

OpenStudy (anonymous):

\[\sqrt{\frac{ \pm \sqrt{17}-3 }{ 2 }}\]

OpenStudy (anonymous):

I need to stop underestimating myself. Got it right first time

OpenStudy (lgbasallote):

...wait..why is there a big square root?

OpenStudy (anonymous):

@ArkGoLucky thats bad.....

OpenStudy (anonymous):

@lgbasallote first you z came than the y... lolzzzZZZ...

OpenStudy (anonymous):

oh sorry then x...

OpenStudy (lgbasallote):

i didn't want to look as if i copied @ParthKohli

OpenStudy (lgbasallote):

anyway...what's with the big square root?

OpenStudy (anonymous):